OpenStudy (anonymous):

? Solve for x. 5x - 11 = 7x - 3

Parth (parthkohli):

5x = 7x + 8 2x + 8 = 0 2x = -8 x = -4

OpenStudy (anonymous):

5x - 11 = 7x - 3 5x-7x=-3+11 -2x=8 x=8/-2 x=-4